THE CYCLOPENTADIENYL RADICAL REVISITED: THE EFFECTS OF ASYMMETRIC DEUTERATION OF JAHN-TELLER MOLECULES

\underline{SAMANTHA STROM}, JINJUN LIU, University of Louisville, Department of Chemistry, Louisville, KY 40292.

Asymmetric deuteration of Jahn-Teller active molecules partially lifts the vibronic degeneracy and hence provides a unique approach to understanding the Jahn-Teller effect. Previously, a spectroscopic model was proposed and used to simulate the spectra of the asymmetrically deuterated isotopomers of the methoxy radical.^{a}^{b} The same model has been implemented and successfully simulated the previously reported high-resolution laser-induced fluorescence (LIF) spectra of the asymmetrically deuterated cyclopentadienyl radical (C_{5}H_{4}D and C_{5}HD_{4}).^{c} A joint fitting of the transitions from both of the zero-point levels of the X̃^{2}E′′_{1} ground electronic state, split by the asymmetric deuteration, to the Ã^{2}A′′_{2} state yields one set of molecular constants for both levels, which, when combined with molecular constants of C_{5}H_{5} and C_{5}D_{5}, can be used to determine the molecular geometry and magnitude of the Jahn-Teller distortion. The main goal of this new investigation is to resolve the discrepancy between the experimentally determined and the ab initio calculated Jahn-Teller distortion.^{d} In addition, a theoretic model is proposed to quantitatively reproduce the splitting of the zero-point level due to the asymmetric deuteration and zero-point energies for the Jahn-Teller distorted structures around the conical intersection.

aD. G. Melnik, J. Liu, R. F. Curl, and T. A. Miller, Mol. Phys. 105, 529 (2007).

bD. G. Melnik, J. Liu, M.-W. Chen, T. A. Miller, and R. F. Curl, J. Chem. Phys. 135, 094310 (2011)

cL. Yu, D.W. Cullin, J.M. Williamson, and T.A. Miller, J. Chem. Phys. 98, 2682 (1993).

dM. J. Bearpark, M. A. Robb, and N. Yamamoto, Spectrochim. Acta Part A 55, 639 (1999).